# Computing Certificates of Strictly Positive Polynomials in Archimedean Quadratic Modules

### v1

@@ 3 Extending Averkov’s results
This section discusses two extensions to Averkov’s lemma. Firstly, details about deriving various parameters are given. Secondly, the lemma is generalized to allow arbitrary subsets of $\mathbb{R}^{n}$ for $B$, thus relaxing the requirement that $B$ be compact, as well as making the generalized lemma more useful in wider contexts.
3.1 Estimation of parameters $\epsilon$, $\gamma$, and $N$
The original proof for Lemma 2.3 is semi-constructive: a specific polynomial $h$ in (1) is constructed and proved to satisfy the conditions of the lemma; however, parameters like $\epsilon,\gamma$ and $N$ are existentially introduced without specifying proper values. We provide details on the computation of the values of these parameters in the proposition below so that with these values, the constructed polynomial $h$ in (1) satisfies the conditions of the lemma.
@@ 3.1 Estimation of parameters $\epsilon$, $\gamma$, and $N$
þ Let $g_{1},\ldots,g_{s},f$, and $B$ be as stated in Lemma 2.3 and the parameters $\gamma$, $\epsilon$, and $N$ be such that
$\begin{split}\gamma&=\frac{1}{2}\max_{1\leq i\leq s}\{\max g_{i}(\bm{x})\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{x}\in B\},\\ \epsilon&<\frac{1}{2}\left|\max_{\bm{x}}\{\min_{1\leq i\leq s}g_{i}(\bm{x})\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{x}\in f^{-1}((-\infty,0))\cap B\}\right|,\\ N&>\frac{1}{2}\max\left\{\frac{\log(\mu)-\log(2s\gamma)}{\log(\gamma)-\log(\gamma+\epsilon)},\frac{\log(M+\mu)-\log(2\epsilon)}{\log(\gamma+2\epsilon)-\log(\gamma+\epsilon)}\right\},\end{split}$ (2)
where $M=-\min_{\bm{y}\in B}f(\bm{y})$. Then the polynomial $h$ in (1) satisfies $f-h>0$ on $B$.
@@
From the proof of Lemma 2.3, one can see that $\gamma$ needs to satisfy the condition that $2\gamma\geq g_{i}(\bar{\bm{x}})$ for each $i=1,\ldots,s$ and any $\bar{\bm{x}}\in B$, and thus the value of $\gamma$ in (2) suffices.
The condition for $\epsilon$ is that $a(T)\cap[-2\epsilon,2\gamma]^{s}=\emptyset$. Then we claim that for any $\bar{\bm{x}}\in T$, there exists $g_{i}\leavevmode\nobreak\ (1\leq i\leq s)$ such that $g_{i}(\bar{\bm{x}})<-2\epsilon$. Otherwise, assume that there exists $\bar{\bm{x}}\in T$ such that $g_{i}(\bar{\bm{x}})\geq-2\epsilon$ for $i=1,\ldots,s$, then we have $(g_{1}(\bar{\bm{x}}),\ldots,g_{s}(\bar{\bm{x}}))\in a(T)\cap[-2\epsilon,2\gamma]^{s}$: a contradiction. From this we know that $-2\epsilon>\min_{i=1,\ldots,s}g_{i}(\bm{x})$ for all $\bm{x}\in T$, and thus the value of $\epsilon$ in (2) suffices.
In the end, we prove the inequality for $N$. Let $M=-\min_{\bm{y}\in B}f({\bm{y}})$. Then from the proof of Lemma 2.3, one can see that if both $\mu-sc(N)$ and $M-sc(N)+C(N)$ are positive, then $f-h$ is also positive. From $\mu-sc(N)>0$, it follows that $c(N)=2\gamma\left(\frac{\gamma}{\gamma+\epsilon}\right)^{2N}\!<\!\frac{\mu}{s}$. Therefore, we can set $2N>\frac{\log(\mu)-\log(2s\gamma)}{\log(\gamma)-\log(\gamma+\epsilon)}$. Then with $c(N)<\frac{\mu}{m}$, we can set $C(N)>M+\mu$ so that $-M-sc(N)+M+\mu>0$. From this constraint, we obtain the following inequality for $N$: $2N>\frac{\log(M+\mu)-\log(2\epsilon)}{\log(\gamma+2\epsilon)-\log(\gamma+\epsilon)}$. Taking the maximum of both values in the right hands of the two inequalities above, we have the value for $N$ in (2). ∎
Plugging the detailed values of the parameters in Proposition 3.3 into Averkov’s method (Lemma 2.3), Algorithm 1 below is derived, making Averkov’s result constructive.
Input: A polynomial set $G=\{g_{1},\dots,g_{s}\}\subseteq\mathbb{R}[\bm{x}]$, a polynomial $f\in\mathbb{R}[\bm{x}]$ such that $f>0$ over $\operatorname{S}(G)$, and a compact set $B\subseteq\mathbb{R}^{n}$
Output: A sequence of sum-of-squares multipliers $(\sigma_{1},\ldots,\sigma_{s})$ in $\mathbb{R}[\bm{x}]$ such that $f-\sum_{i=1}^{s}{\sigma_{i}g_{i}}>0$ over $B$
@@ 10       Choose $\gamma$, $\epsilon$, and $N$ as in (2);
We illustrate Algorithm 1 with the polynomial set $G=\{g_{1},g_{2},g_{3}\}\subset\mathbb{R}[x]$, where $g_{1}=(x+4)(x+1)(x-1)(x-4)$, $g_{2}=x+3$, and $g_{3}=-(x-3)$, and a polynomial $f=-(x+2)(x-2)$ which is positive on $\operatorname{S}(G)=[-1,1]$. Consider the compact set $B=[-5,5]$. First, we compute the minimum of $f$ over $B$, which is $M=-21$. Since it is not positive, we follow the else block in the algorithm. The values of $\epsilon$, $\gamma$, and $N$ computed with (2) are $0.446261$, $2.94026$, and $655$ respectively. One can check that the polynomial $f-\sum_{i=1}^{3}{\sigma_{i}g_{i}}$ is positive over $B$ with the returned certificates $\sigma_{i}=\left(\frac{g_{i}-\gamma}{\gamma+\epsilon}\right)^{2N}$ for $i=1,\dots,3$.
@@ 3.2 Generalization of Averkov’s Lemma
This subsection discusses how Averkov’s lemma 2.3 can be extended by generalizing the compact set $B$ to any subset in $\mathbb{R}^{n}$ when the set $G$ has a single generator. Using this result, by setting $B=\mathbb{R}$ in the univariate case, the requirement to add an Archimedean polynomial (whose certificates cannot be easily computed) to the set of generators can be relaxed, thus allowing the computation of the certificates directly with respect to the original generators $G$.
@@
Let $g\in\mathbb{R}[\bm{x}]$ be a polynomial with bounded $\operatorname{S}(g)$ and $f\in\mathbb{R}[\bm{x}]$ be strictly positive on $\operatorname{S}(g)$. Then, for any subset $B$ of $\mathbb{R}^{n}$, there exists a polynomial $h\in\operatorname{QM}(g)$ such that $f-h>0$ over B.
@@
Choose $M>0$ such that $\operatorname{S}(g)\cap\{\bm{x}\in\mathbb{R}^{n}\leavevmode\nobreak\ |\leavevmode\nobreak\ \|\bm{x}\|\geq M\}=\emptyset$. This is always possible because $\operatorname{S}(g)$ is bounded. Let $K=-\frac{1}{2}\sup\{g(\bm{x})\leavevmode\nobreak\ |\leavevmode\nobreak\ \|\bm{x}\|=2M\}>0$. We claim that $\operatorname{S}(g+K)$ is also bounded. Suppose that otherwise $\operatorname{S}(g+K)$ is unbounded. Then by Theorem 2.4.4 in [4], the semialgebraic set $\operatorname{S}(g+K)$ has a finite number of connected components and thus an unbounded connected component. Consequently, there is a line $r:\mathbb{R}\to\mathbb{R}^{n}$ passing through this unbounded component such that the non-negative set $\{t\in\mathbb{R}\leavevmode\nobreak\ |\leavevmode\nobreak\ g(r(t))+K\geq 0\}$ of the univariate polynomial $g(r(t))+K$ is unbounded. If $g(r(t))$ is a constant function, then $g(r(t))+K<\sup\{g(\bm{x})\leavevmode\nobreak\ |\leavevmode\nobreak\ \|\bm{x}\|=2M\}+K=-K<0$. This contradiction implies that $g(r(t))$ is not a constant function. Since $\operatorname{S}(g)$ is bounded, it follows that $g(r(t))\to-\infty$ as $t\to\infty$, which contradicts the assumption.
Now let $\tilde{g}:=\|\bm{x}\|^{2}g=(\sum_{i=1}^{n}x_{i}^{2})g\in\operatorname{QM}(g)$. Then clearly $\operatorname{S}(\tilde{g})$ is bounded. If the set $B$ is bounded, then the condition falls into the case of Lemma 2.3, and thus it suffices to consider the case when $B$ is unbounded. Let $T=\{\bar{\bm{x}}\in B\leavevmode\nobreak\ |\leavevmode\nobreak\ f(\bar{\bm{x}})\leq 0\}$ and $a$ be the function sending $\bm{x}$ to $\tilde{g}(\bm{x})$ from $\mathbb{R}^{n}$ to $\mathbb{R}$. Since $\operatorname{S}(\tilde{g})$ is bounded, there exists $\gamma>0$ such that $\tilde{g}(B)\subseteq(-\infty,2\gamma]$. If there exists $\bar{\bm{x}}\in T$ such that $\tilde{g}(\bar{\bm{x}})\in[0,2\gamma]$, then $\tilde{g}(\bar{\bm{x}})\geq 0$. In other words, $\bar{\bm{x}}\in\operatorname{S}(\tilde{g})$ but $f(\bar{\bm{x}})<0$, which contradicts the fact that $f>0$ on $\operatorname{S}(\tilde{g})$. Assign $\epsilon<\frac{1}{2}\left|\max_{x}\{\tilde{g}(\bar{\bm{x}})\leavevmode\nobreak\ |\leavevmode\nobreak\ \bar{\bm{x}}\in T\}\right|$. Then $\tilde{g}(\bar{\bm{x}})<-2\epsilon$ for $\bar{\bm{x}}\in T$, and thus $\tilde{g}(T)\cap[-2\epsilon,2\gamma]=\emptyset$.
@@
Consider the univariate polynomial $H(t):=t(\frac{t-\gamma}{\gamma+\epsilon})^{2N}\in\mathbb{R}[t]$, where $N\in\mathbb{N}$ is to be fixed below, and let $c(N)$ and $C(N)$ be defined as in the proof of Lemma 2.3. Then we have that $c(N)\geq H(t)\geq 0$ on $[0,2\gamma]$ and $C(N)\leq-H(t)$ on $(-\infty,-2\epsilon)$. Define $h(\bm{x}):=H(\tilde{g}(\bm{x}))$. Then $h=(\sum_{i=1}^{n}x_{i}^{2})(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N}g\in\operatorname{QM}(g)$. Next, we prove that $f-h$ is strictly positive on $B$ with a proper choice of $N$. Note that when $\bm{x}\notin\operatorname{S}(g+K)$ and $\|\bm{x}\|\rightarrow\infty$, with $\tilde{g}(\bm{x})<-K\|\bm{x}\|^{2}$ we know that $\tilde{g}(\bm{x})$ tends to $\infty$- ∞. Firstly we assign a real number $A>0$ such that the compact set $D:=\{\bar{\bm{x}}\in\mathbb{R}^{n}\leavevmode\nobreak\ |\leavevmode\nobreak\ \|\bar{\bm{x}}\|\leq A\}\subset\mathbb{R}^{n}$ satisfies that $\operatorname{S}(g+K)\subset D$ and $\tilde{g}(\bar{\bm{x}})<-K\|\bar{\bm{x}}\|^{2}<-\gamma-2\epsilon$ for any $\bar{\bm{x}}\in\overline{D}$, where $\overline{D}$ stands for the complement of $D$ in $\mathbb{R}^{n}$.
@@
(1) For any $\bar{\bm{x}}\in B\cap D$, we use a similar approach to deal with it as in the proof of Lemma 2.3. If $\tilde{g}(\bar{\bm{x}})\geq-2\epsilon$, then we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq\mu-H(\tilde{g}(\bar{\bm{x}}))\geq\mu-c(N)$, where $\mu$ is as defined in the proof of Lemma 2.3. Otherwise $\tilde{g}(\bar{\bm{x}})<-2\epsilon$, and we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq f(\bar{\bm{x}})+C(N)$. Since $B\cap D$ is compact, $f$ has a lower bound $L:=\min_{\bm{y}\in B\cap D}f(\bm{y})$, and we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq L+C(N)$. Since $c(N)\rightarrow 0$ and $C(N)\rightarrow+\infty$ as $N\rightarrow+\infty$, we know that $f(\bar{\bm{x}})-h(\bar{\bm{x}})>0$ for any $\bar{\bm{x}}\in D$ when
@@
This conclusion follows from the proof of Proposition 3.1.
@@
$\begin{split}f-h&=f-\tilde{g}\left(\frac{\tilde{g}-\gamma}{\gamma+\epsilon}\right)^{2N}=\tilde{g}^{2k}\left(\frac{f}{\tilde{g}^{2k}}-\tilde{g}^{1-2k}(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N}\right)\\ &=\tilde{g}^{2k}\left(\frac{f}{\tilde{g}^{2k}}-\tilde{g}\left(\frac{1-\frac{\gamma}{\tilde{g}}}{\gamma+\epsilon}\right)^{2k}(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N-2k}\right).\end{split}$
@@
$\begin{split}\left|\frac{f}{\tilde{g}^{2k}}\right|&\leq\frac{\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}\bm{x}^{\bm{\alpha}}|}{K^{2k}\|\bm{x}\|^{2k}}\leq\frac{\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}|\cdot\|\bm{x}\|^{|\bm{\alpha}|}}{K^{2k}\|\bm{x}\|^{2k}}\\ &=\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{K^{2k}}\cdot\|\bm{x}\|^{|\bm{\alpha}|-2k}\leq\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{K^{2k}}\cdot A^{|\bm{\alpha}|-2k}\end{split}$
for any $\bm{x}\notin D$ when $2k\geq\deg(f)$. Denote $\Lambda:=\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{K^{2k}}\cdot A^{|\bm{\alpha}|-2k}$. With $\tilde{g}<-\gamma-2\epsilon$, we have

### v2

@@ 3 Extending Averkov’s results
This section discusses two extensions of Averkov’s lemma. Firstly, details about deriving various parameters are given. Secondly, the lemma is generalized to allow arbitrary subsets of $\mathbb{R}^{n}$ for $B$, thus relaxing the requirement that $B$ be compact. These extensions make the generalized lemma to be applicable in wider contexts.
3.1 Estimation of parameters $\epsilon$, $\gamma$, and NN
The original proof for ?THM? LABEL:lemma:c is semi-constructive: a specific polynomial $h$ in (1) is constructed and proved to satisfy the conditions of the lemma; however, parameters like $\epsilon,\gamma$ and $N$ are existentially introduced without specifying proper values. We provide details on the computation of the values of these parameters in the proposition below so that with these values, the constructed polynomial $h$ in (1) satisfies the conditions of the lemma.
@@ 3.1 Estimation of parameters $\epsilon$, $\gamma$, and NN
Let $g_{1},\ldots,g_{s},f$, and $B$ be as stated in ?THM? LABEL:lemma:c and the parameters $\gamma$, $\epsilon$, and $N$ be such that
$\begin{split}\gamma&>\frac{1}{2}\max_{1\leq i\leq s}\max_{\bar{\bm{x}}\in B}g_{i}(\bar{\bm{x}}),\\ \epsilon&<\frac{1}{2}\left|\max_{\bar{\bm{x}}\in f^{-1}((-\infty,0])\cap B}\min_{1\leq i\leq s}g_{i}(\bar{\bm{x}})\right|,\\ N&>\frac{1}{2}\max\left\{\frac{\log(\mu)-\log(2s\gamma)}{\log(\gamma)-\log(\gamma+\epsilon)},\frac{\log(M+\mu)-\log(2\epsilon)}{\log(\gamma+2\epsilon)-\log(\gamma+\epsilon)}\right\},\end{split}$ (2)
where $M=-\min_{\bar{\bm{x}}\in B}f(\bar{\bm{x}})$. Then the polynomial $h$ in (1) satisfies fh>0f-h>0 on $B$.
@@
From the proof of ?THM? LABEL:lemma:c, one can see that $\gamma$ needs to satisfy the condition that $2\gamma\geq g_{i}(\bar{\bm{x}})$ for each $i=1,\ldots,s$ and any $\bar{\bm{x}}\in B$, and thus the value of $\gamma$ in (2) suffices.
The condition for $\epsilon$ is that $a(T)\cap[-2\epsilon,2\gamma]^{s}=\emptyset$. Then we claim that for any $\bar{\bm{x}}\in T$, there exists $g_{i}\penalty 10000\ (1\leq i\leq s)$ such that $g_{i}(\bar{\bm{x}})<-2\epsilon$. Otherwise, assume that there exists $\bar{\bm{x}}\in T$ such that $g_{i}(\bar{\bm{x}})\geq-2\epsilon$ for $i=1,\ldots,s$, then we have $(g_{1}(\bar{\bm{x}}),\ldots,g_{s}(\bar{\bm{x}}))\in a(T)\cap[-2\epsilon,2\gamma]^{s}$: a contradiction. From this we know that $2\epsilon>\min_{i=1,\ldots,s}g_{i}(\bar{\bm{x}})$ for all $\bar{\bm{x}}\in T$, and thus the value of $\epsilon$ in (2) suffices.
In the end, we prove the inequality for $N$. Let $M=-\min_{\bar{\bm{x}}\in B}f(\bar{\bm{x}})$. Then from the proof of ?THM? LABEL:lemma:c, one can see that if both $\mu-sc(N)$ and Msc(N)+C(N)-M-sc(N)+C(N) are positive, then fhf-h is also positive. From $\mu-sc(N)>0$, it follows that $c(N)=2\gamma\left(\frac{\gamma}{\gamma+\epsilon}\right)^{2N}\!<\!\frac{\mu}{s}$. Therefore, we can set $2N>\frac{\log(\mu)-\log(2s\gamma)}{\log(\gamma)-\log(\gamma+\epsilon)}$. Then with $c(N)<\frac{\mu}{m}$, we can set $C(N)>M+\mu$ so that $M-sc(N)+M+\mu>0$. From this constraint, we obtain the following inequality $N>\frac{1}{2}\frac{\log(M+\mu)-\log(2\epsilon)}{\log(\gamma+2\epsilon)-\log(\gamma+\epsilon)}$. Taking the maximum of both values on the right hand side of the two inequalities above, we have the value for $N$ in (2). ∎
Plugging the detailed values of the parameters in ?THM? LABEL:lemma:d into Averko-v’s method (?THM? LABEL:lemma:c), Algorithm 1 below is derived, making Averkov’s result constructive.
Input: A polynomial set $G=\{g_{1},\dots,g_{s}\}\subseteq\mathbb{K}[X]$, a polynomial $f\in\mathbb{K}[X]$ such that $f>0$ over $\operatorname{S}(G)$, and a compact set $B\subseteq\mathbb{R}^{n}$
Output: A sequence of sum-of-squares multipliers $(\sigma_{1},\ldots,\sigma_{s})$ in $\mathbb{K}[X]$ such that $f-\sum_{i=1}^{s}{\sigma_{i}\cdot g_{i}}>0$ over $B$
@@ 8   Choose $\gamma$, $\epsilon$, and $N$ as in (2);
For actual computation with a compact set $B$ in Proposition LABEL:inequality-proof-averkov, one can further require that $B$ is a semi-algebraic set represented by polynomial inequalities or is covered by such a semi-algebraic set. In this way, since the parameters in (1) do not need to be computed exactly, their estimation reduces to a standard optimization problem. For example, the optimal value of the right-hand side of the second inequality in (2) can be formulated as
maximize $t$ (3)
subject to $g_{i}(\bar{\bm{x}})\geq t,\quad i=1,\ldots,s,$
$f(\bar{\bm{x}})\leq 0,\quad\bar{\bm{x}}\in B.$
The problem in (3) can be solved by methods like semidefinite programming (SDP) or cylindrical algebraic decomposition (CAD), for example using numerical solvers such as Mosek ApS (2025) after converting the input into a standard form. Taking the case when $s=2$, since $\min\{g_{1},g_{2}\}=\frac{g_{1}+g_{2}}{2}-|\frac{g_{1}-g_{2}}{2}|$, we add the constraints as $p^{2}=\frac{(g_{1}-g_{2})^{2}}{2}$ and $p>0$. Then the problem can be restated as
maximize $t$
subject to $\frac{g_{1}+g_{2}}{2}-p\geq t,\quad p^{2}=\frac{(g_{1}-g_{2})^{2}}{2},$
$p(\bar{\bm{x}})>0,f(\bar{\bm{x}})\leq 0\mbox{ for any }\bar{\bm{x}}\in B.$
This technique can then be applied recursively for $s>2$.
[…]
@@
We illustrate Algorithm 1 with the example in Scheiderer (2024, 5.5.11). Let polynomial set $G=\{g_{1},g_{2},$ $g_{3},g_{4}\}\subseteq\mathbb{Q}[x_{1},x_{2},x_{3}]$, where $g_{i}=2x_{i}-1(i=1,\ldots,3),g_{4}=1-x_{1}x_{2}x_{3}$. The set $\operatorname{S}(G)\subseteq[\frac{1}{2},4]^{3}$ is compact and a polynomial $f=-(x_{1}+3)(x_{1}-5)-(x_{2}+3)(x_{2}-5)-(x_{3}+3)(x_{3}-5)$ which is positive on $\operatorname{S}(G)$. Then set $B=[-4,4]^{3}$. First, we compute the minimum of $f$ over $B$, which is 24-24. Since it is not positive, we follow the else block in the algorithm. The values of $\epsilon,\gamma$ and $N$ computed with (2) are $3$, 3636, and $21$ respectively.
As noted in Scheiderer (2024, 5.5.11), the quadratic module in the above example is non-Archimedean. This example illustrates that Averkov’s approach does not rely on the Archimedean condition, and, in fact, does not even require the bounded semialgebraic set $\operatorname{S}(G)$. In contrast, the method we present later extends the set $B$ to the entire space and makes essential use of the Archimedean condition. This highlights the critical role that Archimedean property plays in Putinar’s Positivstellensatz.
Example 3.3.
Now we consider the first example in Schmüdgen (2025, Lemma 18). Let $G=\{g_{1},g_{2},g_{3},g_{4}\}\subseteq\mathbb{Q}[x_{1},x_{2}]$, where $g_{1}=x_{1}$, $g_{2}=x_{2}$, $g_{3}=(1-x_{1})(1-x_{2})$, and $g_{4}=2-(x_{1}^{2}-x_{2}^{2})$. The set $\operatorname{S}(G)=[0,1]^{2}$ is compact and a polynomial $f=(x_{1}+1)(2-x_{1})+(x_{2}+1)(2-x_{2})$ which is positive on $\operatorname{S}(G)$. Then set $B=\operatorname{S}(g_{4})=\{\|\bar{\bm{x}}\|\leq\sqrt{2}\}$ is compact. First, we compute the minimum of $f$ over $B$, which is $\frac{5}{2}$. Since it is not positive, we follow the else block in the algorithm. The values of $\epsilon,\gamma$ and $N$ computed with (2) are $\frac{1}{3}$, $4$, and $19$ respectively.
@@ 3.2 Generalization of Averkov’s Lemma
This subsection discusses an extension of Averkov’s ?THM? LABEL:lemma:c by generalizing the compact set $B$ to any subset in $\mathbb{R}^{n}$ when the set $G$ consists of a single generator. Using this result, by setting $B=\mathbb{R}$ in the univariate case, the requirement to add an Archimedean polynomial (whose certificates cannot be easily computed) to the set of generators can be relaxed, thus allowing the computation of the certificates directly with respect to the original generators $G$. Before establishing our result, we first recall the Łojasiewicz inequality at infinity for the polynomial with a compact zero set presented in Gwoździewicz (1998), whose proof employs techniques from algebraic geometry and differential topology.
Theorem 3.4 (Gwoździewicz (1998, Theorem 1)).
Let $F\in\mathbb{R}[X]$ be a polynomial of degree $d>2$ such that the set $F^{-1}(0)$ is compact. Then there exist constants $c,R>0$ such that
$|\bar{\bm{x}}^{(d-1)^{n}-d}||F(\bar{\bm{x}})|\geq c\text{\qquad for all }|\bar{\bm{x}}|>R,$ (4)
where $|\bar{\bm{x}}|=\max\{|x_{1}|,\ldots,|x_{n}|\}$ is the $L_{\infty}$ norm of $\bar{\bm{x}}$.
The author of Gwoździewicz (1998) points out that no known example achieves the exponent $(d-1)^{n}-d$ in ?THM? LABEL:thm:coercive. In most cases, the inequality $|\bar{\bm{x}}^{t}|\cdot|F(\bar{\bm{x}})|\geq c\text{ for all }|\bar{\bm{x}}|>R$ holds for some degree $t<(d-1)^{n}-d$. This suggests that the degree could potentially be sharpened.
@@
Let $g\in\mathbb{R}[X]$ be a polynomial with bounded $\operatorname{S}(g)$ and $f\in\mathbb{R}[X]$ be strictly positive on $\operatorname{S}(g)$. Then, for any subset $B$ of $\mathbb{R}^{n}$, there exists a polynomial $h\in\operatorname{QM}(g)$ such that fh>0f-h>0 over $B$.
@@
The following proof is constructive and consists of three main steps. First, we construct a modified polynomial $\tilde{g}\in\operatorname{QM}(g)$ such that $\operatorname{S}(\tilde{g})=\operatorname{S}(g)$ and $\tilde{g}(\bar{\bm{x}})\to$-$\infty$ uniformly as $\|\bar{\bm{x}}\|\to\infty$. In this way, this new polynomial $\tilde{g}$ thereby exhibits behaviors similar to the Archimedean polynomial. Second, we construct a univariate polynomial $H(t)$ with controlled growth and define $h:=H(\tilde{g})\in\operatorname{QM}(g)$. Third, we show that fhf-h is strictly positive over $B$ by analyzing its behaviors separately on a compact subset and its complement. The former compact case is treated similarly to ?THM? LABEL:lemma:c, while the non-compact case relies on the decay behavior of $\tilde{g}$ at infinity to ensure positivity.
Step 1. Since the non-negative set $\operatorname{S}(g)$ of a polynomial $g$ is bounded, we know that the zero set $g^{-1}(0)\subseteq\operatorname{S}(g)$ is both closed and bounded, and thus compact. Without loss of generality, we assume that $\deg(g)>2$; otherwise, we can use $(1+\|X\|^{2})\cdot g$ whose degree is greater than $2$, its zero set is also compact, and a member of $\operatorname{QM}((1+\|X\|^{2})\cdot g)$ is also a member of $\operatorname{QM}(g)$. By ?THM? LABEL:thm:coercive there exist two constants $c_{1},R_{1}>0$ such that $|g(\bar{\bm{x}})|\geq c_{1}|\bar{\bm{x}}|^{d-(d-1)^{n}}$ for any $\bar{\bm{x}}$ with $|\bar{\bm{x}}|>R_{1}$. It is known that all norms defined on a finite dimensional normed linear space are equivalent. Then, there exists $c_{2}>0$ such that $|\bar{\bm{x}}|\geq c_{2}\|\bar{\bm{x}}\|$, where $\|\bar{\bm{x}}\|=\sqrt{x_{1}^{2}+\cdots+x_{n}^{2}}$. Meanwhile, with $\operatorname{S}(g)$ bounded, there exists $R_{2}>0$ such that $|g(\bar{\bm{x}})|=-g(\bar{\bm{x}})$ for $\|\bar{\bm{x}}\|\geq R_{2}$. Letting $R=\max\{\frac{R_{1}}{c_{2}},R_{2}\}$ and $c=c_{1}c_{2}^{d-(d-1)^{n}}$, we have
$g(\bar{\bm{x}})=|g(\bar{\bm{x}})|\geq c_{1}|\bar{\bm{x}}|^{d-(d-1)^{n}}\geq c\|\bar{\bm{x}}\|^{d-(d-1)^{n}}$
for all $\|\bar{\bm{x}}\|>R$ since $|\bar{\bm{x}}|\geq c_{2}\|\bar{\bm{x}}\|>c_{2}R\geq c_{2}\frac{R_{1}}{c_{2}}$.
Now, let $\tilde{g}=\|X\|^{2\lfloor\frac{(d-1)^{n}}{2}\rfloor+2}\cdot g\in\operatorname{QM}(g)$. For $\|\bar{\bm{x}}\|>R$, we have
$\tilde{g}(\bar{\bm{x}})\leq-c\|\bar{\bm{x}}\|^{d}.$ (5)
[…]
@@
Step 3. Now we prove that fhf-h is strictly positive on $B$ with a proper choice of $N$. Firstly, we assign a real number $A>R>0$ such that the compact set $D:=\{\bar{\bm{x}}\in\mathbb{R}^{n}\penalty 10000\ |\penalty 10000\ \|\bar{\bm{x}}\|\leq A\}\subset\mathbb{R}^{n}$ satisfies that $\operatorname{S}(g)\subset D$ and $\tilde{g}(\bar{\bm{x}})<-c\|\bar{\bm{x}}\|^{d}<-\gamma-2\epsilon$ for any $\bar{\bm{x}}\in\overline{D}$, where $\overline{D}$ stands for the complement of $D$ in $\mathbb{R}^{n}$.
@@
(1) For any $\bar{\bm{x}}\in B\cap D$, we use a similar approach to deal with it as in the proof of ?THM? LABEL:lemma:c. If $\tilde{g}(\bar{\bm{x}})\geq-2\epsilon$, then we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq\mu-H(\tilde{g}(\bar{\bm{x}}))\geq\mu-c(N)$, where $\mu$ is as defined in the proof of ?THM? LABEL:lemma:c. Otherwise, $\tilde{g}(\bar{\bm{x}})<-2\epsilon$, and we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq f(\bar{\bm{x}})+C(N)$. Since $B\cap D$ is compact, $f$ has a lower bound $L:=\min_{\bar{\bm{x}}\in B\cap D}f(\bar{\bm{x}})$, and we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq L+C(N)$. Since $c(N)\rightarrow 0$ and $C(N)\rightarrow+\infty$ as $N\rightarrow+\infty$, we know that $f(\bar{\bm{x}})-h(\bar{\bm{x}})>0$ for any $\bar{\bm{x}}\in D$ when
@@
This conclusion follows from the proof of ?THM? LABEL:inequality-proof-averkov.
@@
$\begin{split}f-h&=f-\left(\frac{\tilde{g}-\gamma}{\gamma+\epsilon}\right)^{2N}\cdot\tilde{g}=\tilde{g}^{2k}\left(\frac{f}{\tilde{g}^{2k}}-\tilde{g}^{1-2k}(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N}\right)\\ &=\tilde{g}^{2k}\left(\frac{f}{\tilde{g}^{2k}}-\tilde{g}\left(\frac{1-\frac{\gamma}{\tilde{g}}}{\gamma+\epsilon}\right)^{2k}(\frac{\tilde{g}-\gamma}{\gamma+\epsilon})^{2N-2k}\right).\end{split}$
@@
$\begin{split}\left|\frac{f(\bar{\bm{x}})}{\tilde{g}^{2k}(\bar{\bm{x}})}\right|&\leq\frac{\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}\bar{\bm{x}}^{\bm{\alpha}}|}{c^{2k}\|\bar{\bm{x}}\|^{2d\cdot k}}\leq\frac{\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}|\cdot\|\bar{\bm{x}}\|^{|\bm{\alpha}|}}{c^{2k}\|\bar{\bm{x}}\|^{2d\cdot k}}\\ &=\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{c^{2k}}\cdot\|\bar{\bm{x}}\|^{|\bm{\alpha}|-2d\cdot k}\leq\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{c^{2k}}\cdot A^{|\bm{\alpha}|-2d\cdot k}\end{split}$
for any $\bar{\bm{x}}\notin D$ when $2d\cdot k\geq\deg(f)$. Denote $\Lambda:=\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{c^{2k}}\cdot A^{|\bm{\alpha}|-2d\cdot k}$. With $\tilde{g}<-\gamma-2\epsilon$, we have